OpenStudy (anonymous):

multiply to simplest form -5/6 divided by -24

OpenStudy (saifoo.khan):

\[-\frac56 \div -24\] \[-\frac 56 \times -\frac{1}{24}\] \[\frac{5 \times 1}{6 \times 24}\]Can you solve now?

OpenStudy (anonymous):

its the first one , how do i solve this problem?

OpenStudy (saifoo.khan):

These are step of the same problem.

OpenStudy (anonymous):

\[-5/6\div-24\]\[=-5\times-24\div6 =-20\]